# Persistent and anti-Persistent Motion in Bounded and Unbounded Space: Resolution of the First-Passage Problem

### v1

@@ I Introduction
Advances in tracking sensors for animals [1], cells [2] and particles [3] have made conspicuous the importance of quantifying the spatio-temporal dynamics of movement. To this end, two fundamental frameworks, the lattice random walk (LRW) [4, 5, 6], and its continuous counterpart, the Brownian walk [7], have been widely employed to model and analyse Markovian trajectories, that is when the dynamics arise from history independent random processes.
@@
One such example of non-Markovian features in spatio-temporal trajectories is persistence or anti-persistence in the movement steps, that is by having, respectively, a higher tendency to continue in the same direction or to reverse it. Although both movement statistics have been observed, persistent motion is more common and has been observed at small scales, e.g., bacteria undergoing run and tumble dynamics [12], as well as being seen at large scales, e.g., animals foraging [13] and pedestrians in dense crowds [14], while anti-persistent motion has been reported in hemocytes [15], cell extracts [16], telomeres of bone osteosarcoma cells [17] through to the evolution of the score difference in a game of basketball [18]. Persistent motion also plays a central role in the exciton coherence question [19, 20].
The ubiquity of persistent motion in many natural and artificial phenomena has generated a wealth of literature on the subject. Within the space-time continuous description much is known about the one-dimensional occupation probability [21, 5, 22, 23, 24]. However, unlike the case of Brownian walks, with correlation in the motion, the standard renewal theory approach [4, 25, 26] to find the first-passage is not sufficient to determine the temporal dependence of transport related quantities such as the first-passage or first-hitting statistics to one or multiple absorbing targets. While advances have been made in bounded space, these studies are limited to one dimension [27, 28] or can only access the mean first-passage time [29].
@@
Despite the long modelling history of persistent motion in the spatially discrete literature, the only major work on the CRW first-passage dynamics has appeared very recently [34]. That study considers an unbounded lattice both in one and two dimensions under the assumption that the walker moves its first step in one predetermined direction. With this simplification the one-dimensional problem may be reduced to a known first-passage quotient [35] and the exact first-passage dynamics to a single target is derived, while in two dimensions the dynamics are obtained asymptotically. Despite this recent progress, determining the general first-passage time dependence to one or multiple targets has remained an elusive task, with insights for the one target case in the bounded space relying exclusively upon studies of the global mean first-passage time in periodically bounded space [36].
In this paper, we generalise the discrete renewal equation and develop a general theory to determine the first-passage probability to a finite number of targets for a CRW in bounded and unbounded domains of arbitrary dimensions. Focusing on bounded space, we first find the analytic representation of the CRW occupation probability when space is bounded by periodic and reflective boundaries in hypercubic lattices and by periodic boundaries in a six-sided hexagonal lattice. We derive a general form of the first-passage probability and elucidate that with persistence, even in one-dimensional domains, the first-passage probability can be multi-modal. Furthermore, we extend the notion of first-passage time to the directional first-passage time, which we define as the probability that the walker reaches a target site from a specific direction for the first time.
@@
The contents of the rest of the paper is laid out as follows. Section II treats the formulation of the unbounded problem. The bounded propagators or Green’s function of the occupation probability in one dimensional domains are derived in Sec. III, while their higher dimensional counterparts, together with the case of the hexagonal lattice, which we use to model chiral peristent motion, are treated in Sec. IV. In Sec. V we turn our attention to the placement of absorbing traps and discuss the relationship between their location and the imposition of absorbing boundary conditions on the lattice. It is here that we generalise the discrete renewal equation to correlated motion. In Sec. VI, we derive the first moments of these distributions, then in Sec. VII, we display the iterative method for calculating transport statistics for cases where the analytic solutions are not known. Concluding remarks and possible future avenues are then discussed in Sec. VIII.
@@ II Unbounded Propagator
where $Q(n,m,t)$ is the occupation probability at time $t$ to be at site $n$ and state $m$. The state (or internal degree of freedom) represents the direction the walker may enter the site, namely a walker in state 1 must have arrived from the left, while a walker in state 2 from the right (see schematic in Fig. 1). The parameter $f$ is the probability of continuing forward in the direction last travelled while $b$ is the probability of backtracking away from the previous direction travelled. $c=1-(f+b)$ is the so-called lazy or sojourn parameter whereby $0<f+b\leq 1$, with $f+b=1$ corresponding to a walker that moves at every time step. When $f>b$ the walker is in the persistent regime, while for $f<b$ the walker is subject to anti-persistence. When $f=b$, one recovers the diffusive dynamics modelled by the LRW. The effect of the lazy parameter is akin to a rescaling of the velocity in the continuous limit of the CRW, the telegrapher’s equation (see supplementary material [41] for details).
@@ III.2 Reflective Barrier
By taking the state level boundary conditions given above, and performing a summation over the internal states, one finds the boundary condition for the entire lattice site. Upon doing so, one find the boundary condition to be $P^{(r)}_{n_{0}}(1,t)=P^{(r)}_{n_{0}}(0,t)$, $P^{(r)}_{n_{0}}(N,t)=P^{(r)}_{n_{0}}(N+1,t)$, which, as expected [47], is the same boundary condition as the diffusive random walker. Since the image set for this boundary condition has been derived recently [6], one can find the reflective propagator as $P^{(r)}_{n_{0}}(n,t)=\sum_{i=1}^{2}\bm{P}^{(r)}_{n_{0}}(n,t)_{i}$ where
@@
We follow ref. [45] and squeeze a torus into a reflecting lattice. More specifically, we transform a periodic domain with $M\in 2\mathbb{Z}$ sites into a reflective one with $N=(M+2)/2$ sites ($N\geq 2$) by summing the probabilities of the sites a distance $l$ on either side of the $1^{\text{st}}$ and $N^{\text{th}}$ site (see Fig. 1b in [45] or Fig. 1 of the present supplementary material [41] for a schematic representation). In doing so, we impose the boundary condition as implemented orignally by Chandrasekhar for the LRW [50], which states that the walker attempting to escape from $N$ is bounced back to $N-1$ with certainty, and similarly for site $n=1$. The reflective propagator with this condition is [45]
@@
$\bm{P}^{(r_{s})}_{n_{0}}(n,t)=\frac{1}{2N-2}$ $\sum_{\kappa=0}^{2N-3}\bigg{[}\exp\left(\frac{-\pi i\kappa(n-n_{0})}{N-1}\right)+\mu(n)\exp\left(\frac{-\pi i\kappa(2N-n-n_{0})}{N-1}\right)\bigg{]}\bm{\lambda}\left(\frac{\pi\kappa}{N-1}\right)^{t}\cdot\bm{U_{\bm{m}_{0}}}.$ (12)
In Fig. 2(a), we plot the reflecting propagator found via both methods (Eqs. (10) and (12)). For times small enough that no boundary interaction has occurred, one cannot distinguish between the two representations, as expected. However, after boundary interaction, the occupation probability differs slightly due to the no-waiting scheme at the boundary that Eq. (12) obeys. Since these sites can be viewed as ‘slippy’, i.e., the chance of leaving the site is higher than the rest of the domain, the steady state distribution shows a bench-like structure, with these boundary sites having a lower probability of being occupied. In fact, it can be shown (see B) that when $f$ and $b$ are not identically one, $P^{(r_{s})}_{n_{0}}(n,t\to\infty)=\frac{1}{(N-1)}$ ($n\neq\{1,N\}$) while $P^{(r_{s})}_{n_{0}}(1,t\to\infty)=P^{(r_{s})}_{n_{0}}(N,t\to\infty)=\frac{1}{2(N-1)}$.
@@ IV Propagators in Higher Dimensions
As mentioned in the introduction, in continuous space modelling persistence in higher dimensions requires the inclusion of an infinite number of internal variables [30]. A discrete representation, on the other hand, requires only $Z$ internal variables, where $Z$ is the coordination number of the lattice. In dimensions larger than one, we require a third parameter, $\ell$, to govern the probability of turning, which we define as a change of movement direction in any direction except turning back. For simplicity, we first assume the probability of moving in any lateral direction, to be uniform, although this can be relaxed with no additional mathematical burden, as we will show below with the hexagonal lattice. Moreover, in the $d$ dimensional hypercubic lattice we demand $f+b+\mathfrak{d}\ell+c^{(\mathfrak{d})}=1$, where $\mathfrak{d}$ is the number of permissible lateral movements, defined as $\mathfrak{d}=Z-2$ and $c^{(\mathfrak{d})}$ is the sojourn probability. The initial condition must also account for the higher number of movement directions, meaning that $\bm{Q}_{\bm{n}_{0}}(\bm{n},0)=\prod_{i=1}^{d}\left(\delta_{n_{i},n_{0_{i}}}\right)\bm{U}_{\bm{m}_{0}}$, where $\bm{U}_{\bm{m}_{0}}=[\alpha_{1},...,\alpha_{2d}]^{\intercal}$.
@@ IV.1 Hypercubic Lattices
The method of images allows us to find the periodic and reflective propagator in arbitrary dimensions. Following the supplementary material [41], we find the propagator in arbitrary dimensions again as a summation of column vector of state probabilities $P^{(\gamma)}_{n_{01},...,n_{0d}}(n_{1},...,n_{d},t)=\sum_{m=1}^{2d}\bm{P}^{(\gamma)}_{n_{01},...,n_{0d}}(n_{1},...,n_{d},t)_{m}$, where
@@ IV.2 Hexagonal Lattice
When considering the hexagonal lattice, it is convenient to use Her’s three axis coordinate system [54] where each coordinate is represented by three linearly dependent coordinates $(n_{1},n_{2},n_{3})$ such that $n_{1}+n_{2}+n_{3}=0$. The finite domain is six-sided, with the number of lattice points $\Omega=3R^{2}+3R+1$ governed by the circumradius of the hexagon $R$. As the coordination number in this case is $Z=6$, we require $f+b+\sum_{j=1}^{4}\ell_{j}+c^{(4)}=1$ and $\bm{U}_{{\bm{m}_{0}}}$ is a $6\times 1$ column vector. Each lateral movement direction is defined in relation to the previous movement step where $\ell_{1}$ governs the probability of a $60^{\circ}$ rotation anticlockwise, $\ell_{2}$ a $120^{\circ}$ anticlockwise rotation, $\ell_{3}$ a $120^{\circ}$ clockwise rotation and $\ell_{4}$ that of a $60^{\circ}$ clockwise rotation (see [41] (Fig. 2) for pictorial representation). With these parameters, we find the propagator for correlated motion in a periodic hexagonal geometry as (see [41])
$\bm{P}^{(\mathcal{H})}_{\bm{n}_{0}}(\bm{n},t)=\Bigg{\{}\frac{\bm{\lambda}^{(\mathcal{H})}(0,0)^{t}}{\Omega}+\frac{1}{\Omega}\sum_{r=0}^{R-1}\sum_{s=0}^{3r+2}e^{\frac{2\pi i\bm{\kappa}\cdot(\bm{n}-\bm{n}_{0})}{\Omega}}\bm{\lambda}^{(\mathcal{H})}\big{(}\bm{\kappa}(r,s)\big{)}^{t}\Bigg{.}+e^{\frac{-2\pi i\bm{\kappa}\cdot(\bm{n}-\bm{n}_{0})}{\Omega}}\bm{\lambda}^{(\mathcal{H})}\big{(}-\bm{\kappa}(r,s)\big{)}^{t}\Bigg{\}}\cdot\bm{U}_{{\bm{m}_{0}}},$ (20)
with the vector $\bm{\kappa}(r,s)=2\pi[\kappa_{1}(r,s),\kappa_{2}(r,s)]/\Omega$ having components $\kappa_{1}(r,s)=R(s+1)+s-r$ and $\kappa_{2}(r,s)=R(2-s+3r)+r+1$ [55] and with $\bm{\lambda}^{(\mathcal{H})}(\cdot,\cdot)$ defined in Eq. (S45) of the supplementary material [41].
@@
Figure 3: (Colour Online). Correlated motion in hexagonal lattices. Panels (a)-(c) show the occupation probability from Eq. (20) at time $t=11$ in a hexagonal domain with circumradius $R=8$ ($\Omega=217$) and initial condition at the origin (shown via the red circles) with $\bm{U_{\bm{m}_{0}}}=\left(1,1,1,1,1,1\right)^{\intercal}/6$. The boundary condition is taken as the right-shift periodic boundary condition, and we provide further details of how this impacts the dynamics in ref. [41]. In panel (a) the walker is ballistic ($f=1$) and there is no steady state. There are instead probability pulses of magnitude $\frac{1}{6}$ that travel around the domain for all time. In panel (b) $f=0.6$, $b=0.06$ and $\ell_{\{1,2,3,4\}}=0.085$, while in panel (c) $f=0.8$, $\ell_{1}=0.2$ ($b=\ell_{\{2,3,4\}}=0$). Panel (d) depicts a sample trajectory, with the same movement probabilities as panel (c). The trajectory is over $2000$ timesteps and the colour of the trajectory indicates the time, with light red indicating short times and dark red indicating long times. The domain is of size $R=700$, which is chosen so the walker does not interact with the boundary, and we omit the boundary for visual ease. The initial condition is shown by the orange circle at (0,0,0) and the coordinate axes for the hexagonal lattice are shown via the black arrows.
An interesting feature of this boundary condition is that the ballistic walker in the periodic hexagonal lattice is guaranteed to visit every site, which we show mathematically in Sec. VI of the supplementary material [41]. In contrast, in the square lattice, either periodic or reflecting, the ballistic walker is confined to a genuine sublattice, i.e. only the coordinates $[(1,n_{0_{2}}),...,(N,n_{0_{2}}),(n_{0_{1}},1),...,(n_{0_{2}},N)]$, which can be evinced as when $b$ and $s$ are identically zero, $\left[\mathbb{I}-\bm{\lambda}\left(\pi\mathcal{N}^{(\gamma)}_{\kappa_{1}},\pi\mathcal{N}^{(\gamma)}_{\kappa_{2}}\right)\right]^{-1}$ is singular away from $\kappa_{1}=\kappa_{2}=0$, which occurs only when the walk is confined to a genuine sublattice [56]. This has interesting implications on the search time of the walkers in these lattices. Clearly, when the target is away from the coordinates of the sublattice, the search time of the walker in the square lattice becomes increasingly longer as persistence is increased [36] until the probability of a first-passage event ever taking place becomes zero. However, on the hexagonal lattice, the ballistic walker will, with probability one, hit the target in $t\leq\Omega$ timesteps making an increase of persistence a good search strategy with the periodic hexagonal boundary condition.
@@
By relaxing the assumption that all lateral movement directions are uniformly distributed, we introduce the possibility of looping patterns into the movement. We present this as a discrete alternative to the model of ‘chiral’ persistent motion [30, 57, 58, 59], where an angular bias is introduced into the turning angle, which is of interest as a model of a charged particle subjected to soft scattering by aligned magnetic domains [30] and particles driven by microscopic torques such as active colloids [60] and spermatozoa [61], and chlamydomonas [62]. While chiral motion appears at first to lend itself to a continuous space representation, it is known that even the two-dimensional case leads to an analytical intractable Fokker-Planck equation [30], meaning occupation probabilities and further transport quantities are difficult to obtain. On the other hand, in the discrete paradigm, to create looping, one merely alters the movement probabilities in $\bm{\lambda}^{(\mathcal{H})}(\cdot,\cdot)$ allowing to obtain with ease the occupation probability.

### v2

@@ I Introduction
Advances in tracking sensors for animals [1], cells [2] and particles [3] have made conspicuous the importance of quantifying the spatio-temporal dynamics of movement. To this end, two fundamental frameworks, the lattice random walk (LRW) [4, 5, 6], and its continuous counterpart, the Brownian walk [7], have been widely employed to model and analyse Markovian trajectories, that is when the dynamics arise from history-independent random processes.
@@
One such example of non-Markovian features in spatio-temporal trajectories is persistence or anti-persistence in the movement steps, that is by having, respectively, a higher tendency to continue in the same direction or to reverse it. Although both movement statistics have been observed, persistent motion, where the particle has inertia and possesses momentum, is more common and has been observed at small scales, e.g., bacteria undergoing run and tumble dynamics [12], as well as being seen at large scales, e.g., animals foraging [13] and pedestrians in dense crowds [14], while anti-persistent motion has been reported in hemocytes [15], cell extracts [16], telomeres of bone osteosarcoma cells [17] through to the evolution of the score difference in a game of basketball [18]. Persistent motion also plays a central role in the exciton coherence question [19, 20].
The ubiquity of persistent motion in many natural and artificial phenomena has generated a wealth of literature on the subject. Within the space-time continuous description much is known about the one-dimensional occupation probability [5, 21, 22, 23, 24]. However, unlike the case of Brownian walks, with correlation in the motion, the standard renewal theory approach [4, 25, 26] is not sufficient to determine the temporal dependence of transport related quantities such as the first-passage or first-hitting statistics to one or multiple absorbing targets. While advances have been made in bounded space, these studies are limited to one dimension [27, 28] or can only access the mean first-passage time [29].
@@
Despite the long modelling history of persistent motion in the spatially discrete literature, the only major work on the CRW first-passage dynamics has appeared very recently [34]. That study considers an unbounded lattice both in one and two dimensions under the assumption that the walker starts by facing in one predetermined direction. With this simplification the one-dimensional problem may be reduced to a known first-passage quotient [35] and the exact first-passage dynamics to a single target is derived, while in two dimensions the dynamics are obtained asymptotically. Despite this recent progress, determining the general first-passage time dependence to one or multiple targets in finite space has remained an elusive task, with insights for the one target case relying exclusively upon studies of the global mean first-passage time in periodically bounded space [36].
In this paper, we generalise the discrete renewal equation and develop a general theory to determine the first-passage probability to a finite number of targets for a CRW in bounded and unbounded domains of arbitrary dimensions. Focusing on bounded space, we first find the analytic representation of the CRW occupation probability when space is bounded by periodic and reflective boundaries in hypercubic lattices and by periodic boundaries in a six-sided hexagonal lattice. By modelling the special properties of the target site as an impurity or defect problem [37, 38, 39, 40] we derive a general form of the first-passage probability and elucidate that with persistence, even in one-dimensional domains, the first-passage probability can be multi-modal. Furthermore, we extend the notion of first-passage time to the directional first-passage time, which we define as the probability that the walker reaches a target site from a specific direction for the first time. Monte Carlo simulations are then used to validate our theoretical findings.
@@
The contents of the rest of the paper is laid out as follows. Section II treats the formulation of the unbounded problem. The bounded propagators or Green’s function of the occupation probability in one dimensional domains are derived in Sec. III, while their higher dimensional counterparts, together with the case of the hexagonal lattice, which we use to model chiral persistent motion, are treated in Sec. IV. In Sec. V we turn our attention to the placement of absorbing traps and discuss the relationship between their location and the imposition of absorbing boundary conditions on the lattice. It is here that we generalise the discrete renewal equation to correlated motion. In Sec. VI, we derive the first moments of these distributions, then in Sec. VII, we display the iterative method for calculating transport statistics for cases where the analytic solutions are not known. Concluding remarks and possible future avenues are then discussed in Sec. VIII.
Figure 1: (Colour Online). A cartoon depicting persistent motion, for simplicity drawn over a one-dimensonal lattice, with each circle representing a lattice site. We consider some entity, displayed here as a bacterium, travelling with momentum. If it is travelling to the right, as shown in the upper track, its momentum will keep it travelling to the right (green arrow) unless a random fluctuation (red arrow) causes it to change direction. The equivalent case is then true with the direction flipped once it starts moving to the left, which is shown on the lower track. The smaller size of the reversing arrows stems from the reduced chance of that occurring compared to continuing to move in the same direction.
@@ II Unbounded Propagator
where $Q(n,m,t)$ is the occupation probability at time $t$ to be at site $n$ and state $m$. The state (or internal degree of freedom) represents the direction the walker may enter the site, namely a walker in state 1 must have arrived from the left, while a walker in state 2 from the right (see schematic in Fig. 2). The parameter $f$ is the probability of continuing forward in the direction last travelled while $b$ is the probability of backtracking away from the previous direction travelled. $c=1-(f+b)$ is the so-called lazy or sojourn parameter whereby $0<f+b\leq 1$, with $f+b=1$ corresponding to a walker that moves at every time step. When $f>b$ the walker is in the persistent regime, shown pictorially in Fig. 1, while for $f<b$ the walker is subject to anti-persistence. When $f=b$, one recovers the diffusive dynamics modelled by the LRW. The effect of the lazy parameter is akin to a rescaling of the velocity in the continuous limit of the CRW, the telegrapher’s equation (see supplementary material [45] for details).
@@ III.2 Reflective Barrier
By taking the state-level boundary conditions given above, and performing a summation over the internal states, one finds the boundary condition for the entire lattice site. Upon doing so, one find the boundary condition to be $P^{(r)}_{n_{0}}(1,t)=P^{(r)}_{n_{0}}(0,t)$, $P^{(r)}_{n_{0}}(N,t)=P^{(r)}_{n_{0}}(N+1,t)$, which, as expected [51], is the same boundary condition as the diffusive random walker. Since the image set for this boundary condition has been derived recently [6], one can find the reflective propagator as $P^{(r)}_{n_{0}}(n,t)=\sum_{i=1}^{2}\bm{P}^{(r)}_{n_{0}}(n,t)_{i}$ where
@@
We follow ref. [49] and squeeze a torus into a reflecting lattice. More specifically, we transform a periodic domain with $M\in 2\mathbb{Z}$ sites into a reflective one with $N=(M+2)/2$ sites ($N\geq 2$) by summing the probabilities of the sites a distance $l$ on either side of the $1^{\text{st}}$ and $N^{\text{th}}$ site (see Fig. 1b in [49] or A for a schematic representation). In doing so, we impose the boundary condition as implemented originally by Chandrasekhar for the LRW [54], which states that the walker attempting to escape from $N$ is bounced back to $N-1$ with certainty, and similarly for site $n=1$. The reflective propagator with this condition is [49]
@@
$\bm{P}^{(r_{s})}_{n_{0}}(n,t)=\frac{1}{2N-2}$ $\sum_{\kappa=0}^{2N-3}\bigg{[}\exp\left(\frac{-\pi i\kappa(n-n_{0})}{N-1}\right)$ (12)
$+\mu(n)\exp\left(\frac{-\pi i\kappa(2N-n-n_{0})}{N-1}\right)\bigg{]}\bm{\lambda}\left(\frac{\pi\kappa}{N-1}\right)^{t}\cdot\bm{U_{\bm{m}_{0}}}.$
In Fig. 3(a), we plot the reflecting propagator found via both methods (Eqs. (10) and (12)). For times small enough that no boundary interaction has occurred, one cannot distinguish between the two representations, as expected. However, after boundary interaction, the occupation probability differs slightly due to the no-waiting scheme at the boundary that Eq. (12) obeys. Since these sites can be viewed as ‘slippy’, i.e., the chance of leaving the site is higher than the rest of the domain, the steady state distribution shows a bench-like structure, with these boundary sites having a lower probability of being occupied. In fact, it can be shown (see B) that when either $f$ or $b$ are not identically one, $P^{(r_{s})}_{n_{0}}(n,t\to\infty)=\frac{1}{(N-1)}$ ($n\neq\{1,N\}$) while $P^{(r_{s})}_{n_{0}}(1,t\to\infty)=P^{(r_{s})}_{n_{0}}(N,t\to\infty)=\frac{1}{2(N-1)}$.
@@ IV Propagators in Higher Dimensions
As mentioned in the introduction, in continuous space modelling persistence in higher dimensions requires the inclusion of an infinite number of internal variables [30]. A discrete representation, on the other hand, requires only $Z$ internal variables, where $Z$ is the coordination number of the lattice. In dimensions larger than one, we require a third parameter, $\ell$, to govern the probability of turning, which we define as a change of movement direction in any direction except turning back. For simplicity, we first assume the probability of moving in any lateral direction, to be uniform, although this can be relaxed with no additional mathematical burden, as we will show below with the hexagonal lattice. Moreover, we demand $f+b+\mathfrak{d}\ell+c^{(\mathfrak{d})}=1$, where $\mathfrak{d}$ is the number of permissible lateral movements, defined as $\mathfrak{d}=Z-2$ and $c^{(\mathfrak{d})}$ is the sojourn probability. The initial condition must also account for the higher number of movement directions, meaning that $\bm{Q}_{\bm{n}_{0}}(\bm{n},0)=\prod_{i=1}^{d}\left(\delta_{n_{i},n_{0_{i}}}\right)\bm{U}_{\bm{m}_{0}}$, where $\bm{U}_{\bm{m}_{0}}=[\alpha_{1},...,\alpha_{2d}]^{\intercal}$.
@@ IV.1 Hypercubic Lattices
The method of images allows us to find the periodic and reflective propagator in arbitrary dimensions. Following the supplementary material [45], we find the propagator in arbitrary dimensions again as a summation of column vector of state probabilities $P^{(\gamma)}_{n_{01},...,n_{0d}}(n_{1},...,n_{d},t)=\sum_{i=1}^{2d}\bm{P}^{(\gamma)}_{n_{01},...,n_{0d}}(n_{1},...,n_{d},t)_{i}$, where
@@ IV.2 Hexagonal Lattice
When considering the hexagonal lattice, it is convenient to use Her’s three axis coordinate system [58] where each coordinate is represented by three linearly dependent coordinates $(n_{1},n_{2},n_{3})$ such that $n_{1}+n_{2}+n_{3}=0$. The finite domain is six-sided, with the number of lattice points $\Omega=3R^{2}+3R+1$ governed by the circumradius of the hexagon $R$. As the coordination number in this case is $Z=6$, we require $f+b+\sum_{j=1}^{4}\ell_{j}+c^{(4)}=1$ and $\bm{U}_{{\bm{m}_{0}}}$ is a $6\times 1$ column vector. Each lateral movement direction is defined in relation to the previous movement step where $\ell_{1}$ governs the probability of a $60^{\circ}$ rotation anticlockwise, $\ell_{2}$ a $120^{\circ}$ anticlockwise rotation, $\ell_{3}$ a $120^{\circ}$ clockwise rotation and $\ell_{4}$ that of a $60^{\circ}$ clockwise rotation, see Fig. 12 (Appendix C) for pictorial representation. With these parameters, we find the propagator for correlated motion in a periodic hexagonal geometry as (see C)
$\bm{P}^{(\mathcal{H})}_{\bm{n}_{0}}(\bm{n},t)=\Bigg{\{}\frac{\bm{\lambda}^{(\mathcal{H})}(0,0)^{t}}{\Omega}+$ $\frac{1}{\Omega}\sum_{r=0}^{R-1}\sum_{s=0}^{3r+2}e^{\frac{2\pi i\bm{\kappa}\cdot(\bm{n}-\bm{n}_{0})}{\Omega}}\bm{\lambda}^{(\mathcal{H})}\big{(}\bm{\kappa}(r,s)\big{)}^{t}\Bigg{.}$
$+e^{\frac{-2\pi i\bm{\kappa}\cdot(\bm{n}-\bm{n}_{0})}{\Omega}}\bm{\lambda}^{(\mathcal{H})}\big{(}-\bm{\kappa}(r,s)\big{)}^{t}\Bigg{\}}\cdot\bm{U}_{{\bm{m}_{0}}},$ (20)
with the vector $\bm{\kappa}(r,s)=2\pi[\kappa_{1}(r,s),\kappa_{2}(r,s)]/\Omega$ having components $\kappa_{1}(r,s)=R(s+1)+s-r$ and $\kappa_{2}(r,s)=R(2-s+3r)+r+1$ [59] and with $\bm{\lambda}^{(\mathcal{H})}(\cdot,\cdot)$ defined in Eq. (38).
@@
Figure 4: (Colour Online). Correlated motion in hexagonal lattices. Panels (a)-(c) show the occupation probability from Eq. (20) at time $t=11$ in a hexagonal domain with circumradius $R=8$ ($\Omega=217$) and initial condition at the origin (shown via the red circles) with $\bm{U_{\bm{m}_{0}}}=\left(1,1,1,1,1,1\right)^{\intercal}/6$. The boundary condition is taken as the right-shift periodic boundary condition, and we provide further details of how this impacts the dynamics in C. In panel (a) the walker is ballistic ($f=1$) and there is no steady state. There are instead probability pulses of magnitude $\frac{1}{6}$ that travel around the domain for all time. In panel (b) $f=0.6$, $b=0.06$ and $\ell_{\{1,2,3,4\}}=0.085$, while in panel (c) $f=0.8$, $\ell_{1}=0.2$ ($b=\ell_{\{2,3,4\}}=0$). Panel (d) depicts a sample trajectory, with the same movement probabilities as panel (c). The trajectory is over $2000$ timesteps and the colour of the trajectory indicates the time, with light red indicating short times and dark red indicating long times. The domain is of size $R=700$, which is chosen so the walker does not interact with the boundary, and we omit the boundary for visual ease. The initial condition is shown by the orange circle at (0,0,0) and the coordinate axes for the hexagonal lattice are shown via the black arrows.
An interesting feature of this boundary condition is that the ballistic walker in the periodic hexagonal lattice is guaranteed to visit every site (see C for mathematical details). In contrast, in the square lattice, either periodic or reflecting, the ballistic walker is confined to a genuine sublattice, i.e. only the coordinates $[(1,n_{0_{2}}),...,(N,n_{0_{2}}),(n_{0_{1}},1),...,(n_{0_{2}},N)]$, which can be evinced as when $b$ and $\ell$ are identically zero, $\left[\mathbb{I}-\bm{\lambda}\left(\pi\mathcal{N}^{(\gamma)}_{\kappa_{1}},\pi\mathcal{N}^{(\gamma)}_{\kappa_{2}}\right)\right]^{-1}$ is singular away from $(\kappa_{1},\kappa_{2})=(0,0)$, which occurs only when the walk is confined to a genuine sublattice [60]. This has interesting implications on the search time of the walkers in these lattices. Clearly, when the target is away from the coordinates of the sublattice, the search time of the walker in the square lattice becomes increasingly longer as persistence is increased [36] until the probability of a first-passage event ever taking place becomes zero. However, on the hexagonal lattice, the ballistic walker will, with probability one, hit the target in $t\leq\Omega$ timesteps making an increase of persistence a good search strategy with the periodic hexagonal boundary condition.
@@
By relaxing the assumption that all lateral movement directions are uniformly distributed, we introduce the possibility of looping patterns into the movement. We present this as a discrete alternative to the model of ‘chiral’ persistent motion [30, 61, 62, 63], where an angular bias is introduced into the turning angle, which is of interest as a model of a charged particle subjected to soft scattering by aligned magnetic domains [30] and particles driven by microscopic torques such as active colloids [64], spermatozoa [65], and chlamydomonas [66]. While chiral motion appears at first to lend itself to a continuous space representation, it is known that even the two-dimensional case leads to an analytical intractable Fokker-Planck equation [30], meaning occupation probabilities and further transport quantities are difficult to obtain. On the other hand, in the discrete paradigm, to create looping, one merely alters the movement probabilities in $\bm{\lambda}^{(\mathcal{H})}(\cdot,\cdot)$ allowing to obtain with ease the occupation probability.